# CruxEval-X: A Benchmark for Multilingual Code Reasoning, Understanding and Execution

### v1

@@ Introduction
Figure 3: The result of each LLM in CruxEval-X. Each result is shaded with a background color from blue to white based on the Pass@1. The bluer, the larger. Deepseekcoder-V2 is a MOE LLM; the parameter activated during inference is 21B.
@@
LLMs for evaluation We selecte 24 LLMs across 4 types for evaluation, including general LLMs (GPT-3.5-Turbo, GPT-4o-mini, GPT-4o (Brown et al. 2020; Achiam et al. 2023), Llama3 (AI 2024), Qwen2 (Yang et al. 2024a), phi-3-instruct (Abdin et al. 2024)), multilingual code LLMs (Deepseekcoder-V2 (Zhu et al. 2024), Deepseekcoder-V1 (Guo et al. 2024), CodeLlama (Roziere et al. 2023), Starcoder (Li et al. 2023a), Starcoder2 (Lozhkov et al. 2024), CodeQwen1.5-Chat (Bai et al. 2023)), instruction-tuned multilingual LLMs (Deepseekcoder-instruct-V1 (Guo et al. 2024), WizardCoder (Luo et al. 2023), CodeLlama-Python, CodeLlama-Instruct (Roziere et al. 2023)), and single or few-language code LLMs (CodeGen (Nijkamp et al. 2022), phi-1 (Gunasekar et al. 2023), phi-1.5 (Li et al. 2023b)).
Evaluation task We adopt the task settings from prior work (Gu et al. 2024), dividing the tasks into output reasoning and input reasoning. For any PLs dataset $L^{k}\in\{L_{i}\}^{K}_{i=1}$, where $K=19$, representing the total number of PLs. We provide a function $f^{L_{k}}$ and the corresponding test. The input of this test example is $i^{L_{k}}$, and the output is $o^{L_{k}}$.
The output reasoning task can be expressed as:
$r^{L_{k}}=I(P(o^{L_{k}}|f^{L_{k}},i^{L_{k}},M))$ (2)
The input reasoning task can be expressed as:
$r^{L_{k}}=I(P(i^{L_{k}}|f^{L_{k}},o^{L_{k}},M))$ (3)
where $M$ is any LLMs. We get the input or output from the code generation distribution $P(.)$ and compose test cases $(i,o)^{L^{k}}$. $I(.)$ is the indicator function by executing this test case with function $f^{L_{k}}$. If $f^{L_{k}}$ passes the test, the evaluation result is 1, otherwise 0.
@@
Evaluation method. We use pass@1 (Kulal et al. 2019; Chen et al. 2021) as the evaluation metric to assess both the task of output reasoning and input reasoning. We set the temperature to 0 and employ greedy decoding for generation as prior work (Cao et al. 2024a). For closed-source LLMs, we generate outputs by calling OpenAI’s API. The version of the api is gpt-3.5-turbo, gpt-4o-mini, gpt-4o
Overall Result.
@@
Figure 3 shows Pass@1 evaluation results of various LLMs arranged in descending order of the LLMs’ parameter size. It can be observed that LLMs with a larger number of parameters tend to perform better. The closed-source LLMs GPT-4o achieved the best results among all evaluated LLMs. Notably, the result of the open-source LLMs Deepseekcoder-V2 is better than GPT-4o-mini, with an average Pass@1 of 62.8% on input reasoning and 65.0% on output reasoning. From the perspective of average values, it can be observed that under different types of PLs, the LLMs’ capabilities on input and output reasoning are quite similar, which echos the observation made by prior work (Gu et al. 2024) in Python. More interestingly, during the evaluation, we introduced the few-language LLMs CodeGen-muti, which was only trained on C, C++, GO, Java, JavaScript, and Python, as well as the single-language LLMs phi-1.5 and phi-1, which were pre-trained only on Python. However, they achieve similar results in the input and output reasoning tasks across 19 PLs. Notably, phi-1, having only seen Python, achieved a Pass@1 of 11.8% on Python input prediction, but scored 23.6% on Perl input prediction. We will provide a more detailed analysis of this phenomenon in Analysis part.
Analysis on Key Factors for LLM Code Reasoning
@@
To get more insight into what factors in code affect LLMs’ code reasoning ability, we explore six factors (e.g., average number of input variables, average input length) and statistics their correlation with correct/incorrect reasoning. The result is shown in Figure 4. Each column of the box plot displays the distribution of 19 PLs. In particular, the columns “Num of Input Variable”, “Num of Variable Type” are counted by averaging the number of types of input parameters in method signatures, “Input/Output Length” is the average string length of the input/output. For instance, in the test case where f("a", 123) == 123, the input length is 4, and the output length is 3. “Whether Have Complex Type” checks whether there are List, Dict, Tuple, Set types in input and output. “Execution Steps” calculates the average execution steps in Python bytecode operations, following prior work (Gu et al. 2024).
@@
From Figure 4, we can see from Sub-figures A-C that the number/types of input variables have little impact on the code reasoning, especially Sub-figure B, which shows that the reasoning capability is slightly better when more types of variables are involved. A more counter-intuitive observation is made from sub-figures D-E. They indicate that the reasoning capability is negatively correlated with the length of input/output strings instead of the number of data types.
@@
Furthermore, regarding input reasoning capability, the more input variables, the more challenging it is for LLMs to reason about the correct inputs, thus the worse the input reasoning performance (Sub-figures A and D). Similarly, the longer the outputs, the harder the output reasoning, resulting in worse performance (Sub-figure E).
@@
Analysis on Cross-language Generalization
To investigate the cross-language generalizability of LLMs, we investigate the reasoning ability of phi-1 and phi-1.5, which are trained on English and Python only. To get a better understanding, we analyze the capability in terms of syntax and semantics in 9 PLs because they provide clear error messages to distinct syntactic/semantic errors.
Syntactic Correctness.
To generalize to other PLs, it is critical to ensure syntactic correctness. Figure 5 (A)-(B) show the number of syntactic-correct cases made by these two LLMs in both tasks. It is clear that Python has the highest syntactic correctness in both tasks, followed by Go and TypeScript. On the contrary, C++, C#, and Java witness the most syntactic errors for three LLMs. Interestingly, even though phi-1 and phi-1.5 have not trained on PLs other than Python, they can still achieve an average of 49.1% and 72.0% syntactically correctness rate in other PLs, respectively, compared with 97.0% and 98.7% achieved on Python. It indicates the cross-language generalizability of LLMs.
Semantic Correctness.
Beyond syntactic correctness, semantic correctness poses higher requirements, i.e., passing the tests. The results are shown in the last two rows in Figure 3. Similar results can be observed in both tasks and both LLMs. In particular, phi-1.5 reaches 25.8% input reasoning performance on Python, while on other PLs, an average of 19.0% can also be reached. The observation further consolidates the cross-language generalizability of LLMs.
Cross-NL and Cross-PL Generalization.
From Figure 5 and Figure 3, there is a noticeable increase from phi-1 to phi-1.5 (an average of 10.7% vs. 21.7% on input reasoning, and 15.1% vs. 21.7% on output reasoning). According to the description (Abdin et al. 2024), phi-1.5 is further fine-tuned with more synthetic texts in natural language (NL). Considering the dramatic improvement in code reasoning, it is highly likely that the improvement in NL reasoning positively impacts code reasoning.
Figure 5: The number of syntax error of each LLM.
Analysis on Programming Language Correlation
@@
Case Study
After identifying the phi-series-LLMs (i.e., phi-1, phi-1.5) exhibit cross-language generalization and the correlation across PLs, we further analyze the predictions of phi-1.5 to get a deeper understanding. We noticed that out of the 128 correct instances in Python by phi-1.5, 61.7% (79/128) are also correct in PHP, while only 39.8% (51/128) are correct in Racket. Therefore, we use one example in these three PLs (Python, PHP, and Racket) to understand its rationale.
Analysis on Subject 106.
Listing 1-3 demonstrates an instance where phi-1.5 generalizes Python (Listing 1)’s reasoning capabilities to other languages. Each example’s check function has been uniformly simplified to assert f() == (). From a grammar structure perspective, their respective function definitions, indentation formats, and the functions they invoke exhibit significant differences. However, overall, PHP and Python share a more similar structure, both utilizing sort for sorting and return for returning output values. Therefore, phi-1.5 is able to generalize its code reasoning abilities to PHP, but fails to comprehend the sorting command in Racket, leading to incorrect predictions.
Upon analyzing these 128 questions, we observe that excluding those where the output could be directly derived from the input, such as assert f("zej","owc") == "zej", which accounted for approximately 40% of the cases, there are still numerous examples demonstrating that phi-1.5 has developed a certain level of cross-language capabilities. From these examples, we can observe that the multilingual generalization capability of the model is positively correlated with the grammar structural similarity between languages. Even Racket, a language significantly different from others, maintains certain logical similarities in aspects such as function definitions, loops, and conditional branchs. This is a key reason why Phi-1.5 can achieve considerable effectiveness across multiple languages.
Listing 1 Subject-106 (Python)
@@
Listing 3 Subject-106 (Racket)
@@
4(require rackunit)
5assert f(list 5 8 1 3 0) == ????
6# phi-1.5 answer: list 5 8 1
Related Work
Multi-Task Code Benchmark Recently, there has been an increasing number of tasks related to code that are used to evaluate the various capabilities of LLMs in the field of coding, including code generation (Chen et al. 2021; Austin et al. 2021), code repair (Jimenez et al. 2023; Tian et al. 2024), and code description (Chai et al. 2024). However, datasets that assess the reasoning abilities of code are relatively limited, and the currently proposed reasoning datasets are confined to the Python language (Gu et al. 2024; Chen et al. 2024). In this work, we have expanded the Python language reasoning dataset CRUXEVAL (Gu et al. 2024) to encompass 19 PLs, thereby addressing the deficiency in reasoning datasets at the multilingual level.
Multi-Language Code Benchmark Multilingual evaluation datasets are an important method for assessing the comprehensive coding capabilities of code LLMs. In the early stages, multilingual code datasets were mainly used for code translation tasks (Elnaggar et al. 2021; Ahmad et al. 2021; Roziere et al. 2020, 2021; Zhu, Suresh, and Reddy 2022; Yan et al. 2023; Zhu et al. 2022). These datasets often consist of problem solutions in different languages extracted from algorithm competition-related websites, thus suffering from data contamination issue. Bechmark like McEval (Chai et al. 2024), which relies on human annotation, requires a high cost. In this work, we provide a process using LLMs for multilingual code translation, which can achieve a high accuracy and low cost in creating a multilingual dataset.
Conclusion
@@
Appendix A Appendix
The improvements of MutiPL-E pipline
enhance of the pipline
@@ Appendix A Appendix
transform complex types
@@
Listing 4 example with complex type
@@
The difficulties for translating
(1) For indexing functions such as ‘index‘, the starting position is not 0 but 1, with Julia being a typical language that exhibits this behavior. (2) For the conversion between Python”s ‘str‘ type and its own ‘char‘ and ‘string‘ types, D language is a typical case where this issue arises. (3) For the transformation and comparison of dictionary types, C# is a typical language where such errors are common.
Prompt of Benchmark Construction
In Figures 7 and 8, we include the prompts we use for our benchmark construction. We use a few-shot prompt for all models. For Generation step of each model, the prompt is show in Figure 7. The example in this Figure is used for GPT3.5-turbo. After the Generation and Repair of GPT3.5-turbo, we choose three examples from the correct generated problems which include str, List, Dict type respectively. We use these examples for Deepseekcoder-instruct-33b Generation. For GPT-4o, based on the summarized difficulties in translation, we provide three examples, as shown in Listing 5, 6, and 7. we present these examples in three distinct languages. For each specific language translation, we employ the corresponding language version of the three examples.
Figure 7: Prompt of Generation
Figure 8: Prompt of Repair

### v2

@@ 3.1 Experiment Setup
Table 2: The result of each LLM in CruxEval-X. Each result is shaded with a background color from blue to white based on the Pass@1. The bluer, the larger. Deepseekcoder-V2 is a MOE LLM; the parameter activated during inference is 21B.
@@
LLMs for evaluation We selected 24 LLMs across 4 types for evaluation, including general LLMs (GPT-3.5-Turbo, GPT-4o-mini, GPT-4o (Brown et al., 2020; Achiam et al., 2023), Llama3 (AI, 2024), Qwen2 (Yang et al., 2024b), phi-3-instruct (Abdin et al., 2024)), multilingual code LLMs (Deepseekcoder-V2 (Zhu et al., 2024), Deepseekcoder-V1 (Guo et al., 2024), CodeLlama (Roziere et al., 2023), Starcoder (Li et al., 2023a), Starcoder2 (Lozhkov et al., 2024), CodeQwen1.5-Chat (Bai et al., 2023)), instruction-tuned multilingual LLMs (Deepseekcoder-instruct-V1 (Guo et al., 2024), WizardCoder (Luo et al., 2023), CodeLlama-Python, CodeLlama-Instruct (Roziere et al., 2023)), and single or few-language code LLMs (CodeGen (Nijkamp et al., 2022), phi-1 (Gunasekar et al., 2023), phi-1.5 (Li et al., 2023b)).
Evaluation task We adopt the task settings from prior work (Gu et al., 2024), dividing the tasks into output reasoning and input reasoning. For any PLs dataset, we provide the code along with the corresponding input or output in the test cases. Input reasoning is predicting the input based on the output. Output reasoning is predicting the output based on the input. Each problem contains a test case, which follows the CRUXEVAL. The search space for test cases is exponential in nature, ensuring that LLMs do not know the answer to the problem in advance.
@@
Evaluation method. We use pass@1 (Kulal et al., 2019; Chen et al., 2021) to evaluate both tasks. We set the temperature to 0 and employ greedy decoding for generation as prior work (Cao et al., 2024a). For closed-source LLMs, we use gpt-3.5-turbo, gpt-4o-mini, and gpt-4o from OpenAI’s API.
3.2 Overall Result
@@
Table 2 shows Pass@1 evaluation results of various LLMs arranged in descending order of the LLMs’ parameter size. We can see that: 1) CruxEval-X is challenging for all tested LLMs. Even GPT-4o can only achieve Pass@1 around 70%. Notably, the result of the open-source LLMs Deepseekcoder-V2 is better than GPT-4o-mini, with an average Pass@1 of 62.8% on input reasoning and 65.0% on output reasoning. Furthermore, the average results show that across different types of PLs, LLMs exhibit similar capabilities in input and output reasoning, consistent with prior observations in Python (Gu et al., 2024). 2) Single-programming-language model (e.g., phi-1) and few-language models (e.g. CodeGen-muti) demonstrate unexpected generalization capabilities on unseen language tasks. During evaluation, we introduced few-language LLMs (CodeGen-multi, trained on six PLs) and single-language LLMs (phi-1.5 and phi-1, trained only on Python). Despite this, they achieved similar results across 19 PLs. Notably, phi-1, trained solely on Python, scored 11.8% Pass@1 on Python input prediction and 23.6% on Perl. We will provide a more detailed analysis of this phenomenon in the Analysis section.
4 Analysis
4.1 Key Factors for LLM Code Reasoning
@@
To get more insight into what factors in code affect LLMs’ code reasoning ability, we explore six factors (e.g., average number of input variables, average input length) and statistics their correlation with correct/incorrect reasoning. The results are shown in Figure 3. Each column of the box plot displays the distribution of 19 PLs. In particular, the columns “Num of Input Variable”, “Num of Variable Type” are counted by averaging the number of types of input parameters in method signatures, “Input/Output Length” is the average string length of the input/output. For instance, in the test case where f(’a’, 123) == {’a’: 1}, the input length is 6, and the output length is 8. “Whether Have Complex Type” checks whether there are List, Dict, Tuple, Set types in input and output. “Execution Steps” calculates the average execution steps in Python bytecode operations, following prior work (Gu et al., 2024).
@@
From Figure 3 ( A - C ), we can see that the number/types of input variables have little impact on the code reasoning, especially Sub-figure B, which shows that the reasoning capability is slightly better when more types of variables are involved. A more counter-intuitive observation is made from Figure 3 ( D - E ). They indicate that the reasoning capability is negatively correlated with the length of input/output strings.
@@
Furthermore, regarding input reasoning capability, the more input variables, the more challenging it is for LLMs to reason about the correct inputs, thus the worse the input reasoning performance (Sub-figures A and D).
@@
4.2 Cross-language Generalization Ability
To investigate the cross-language generalization ability of LLMs, we investigate the reasoning ability of phi-1 and phi-1.5, which are trained on English and Python only. To get a better understanding, we analyze the capability in terms of syntax and semantics in 9 PLs because they provide clear error messages to distinct syntactic/semantic errors.
4.2.1 Syntactic Correctness.
For a LLM trained on single PL, ensuring syntactic correctness is crucial for its generalization to other programming languages. For example, in a Racket language test case, the expected output might be "(list 5 8 1 3 0)", but the LLM incorrectly predicts "[5,8,1,3,0]", leading to a compilation error. Figure 4 shows the number of syntactic-correct cases made by these two LLMs in both tasks. It is clear that Python has the highest syntactic correctness in both tasks, followed by Go and TypeScript. On the contrary, C++, C#, and Java witness the most syntactic errors for these LLMs. Interestingly, even though phi-1 and phi-1.5 have not trained on PLs other than Python, they can still achieve an average of 49.1% and 72.0% syntactically correctness rate in other PLs, respectively, compared with 97.0% and 98.7% achieved on Python. It indicates the cross-language generalizability of LLMs.
4.2.2 Semantic Correctness.
Beyond syntactic correctness, semantic correctness poses higher requirements, i.e., passing the tests. The results are shown in the last two rows in Table 2. In particular, phi-1.5 reaches 25.8% input reasoning performance on Python, while on other PLs, an average of 19.0% can also be reached. The observation further consolidates the cross-language generalizability of LLMs.
4.2.3 Cross-NL and Cross-PL Generalization.
From Figure 4 and Table 2, there is a noticeable increase from phi-1 to phi-1.5 (an average of 10.7% vs. 21.7% on input reasoning, and 15.1% vs. 21.7% on output reasoning). According to the description (Abdin et al., 2024), phi-1.5 is further fine-tuned with more synthetic texts in natural language (NL). Considering the dramatic improvement in code reasoning, it is highly likely that the improvement in NL reasoning positively impacts code reasoning.
Figure 4: The syntax accuracy of each LLM.
4.3 Programming Language Correlation
@@
4.4 Case Study
After identifying the phi-series-LLMs (i.e., phi-1, phi-1.5) exhibit cross-language generalization and the correlation across PLs, we further analyze the predictions of phi-1.5 to get a deeper understanding. We noticed that out of the 128 correct instances in Python by phi-1.5, 61.7% (79/128) are also correct in PHP, while only 39.8% (51/128) are correct in Racket. Therefore, we use one example in these three PLs to understand its rationale.
Listing 1 Subject-164 (Python)
@@ 4.4 Case Study
Listing 3 Subject-164 (Racket)
@@
4;; input: (list 5 8 1 3 0)
5;; phi-1.5 answer: (list 5 8 1)
4.4.1 Analysis on Subject 164.
Listing 1-3 demonstrates an instance where phi-1.5 generalizes Python (Listing 1)’s reasoning capabilities to other languages. From a grammar structure perspective, their respective function definitions, indentation formats, and the functions they invoke exhibit significant differences. However, overall, PHP and Python share a more similar structure, both utilizing sort for sorting and return for returning output values. Therefore, phi-1.5 is able to generalize its code reasoning abilities to PHP, but fails to comprehend the sorting command in Racket, leading to incorrect predictions.
Upon analyzing these 128 questions, we observe that excluding those where the output could be directly derived from the input, such as assert f("zej","owc") == "zej", which accounted for approximately 40% of the cases, there are still numerous examples demonstrating that phi-1.5 has developed a certain level of cross-language capabilities. From these examples, we can observe that the multilingual generalization capability of the model is positively correlated with the grammar structural similarity between languages. Even Racket, a language significantly different from others, maintains certain logical similarities in aspects such as function definitions, loops, and conditional branchs. This is a key reason why Phi-1.5 can achieve considerable effectiveness across multiple languages. More case studies can be found in appendix K.
5 Related Work
Multi-Task Code Benchmark. Recently, there has been an increasing number of tasks related to code that are used to evaluate the various capabilities of LLMs in the field of coding, including code generation (Chen et al., 2021; Austin et al., 2021), code repair (Jimenez et al., 2023; Tian et al., 2024), and code description (Chai et al., 2024). However, datasets that assess the reasoning abilities of code are relatively limited, and the currently proposed reasoning datasets are confined to the Python language (Gu et al., 2024; Chen et al., 2024). In this work, we expand the Python language reasoning dataset CRUXEVAL (Gu et al., 2024) to encompass 19 PLs, thereby addressing the deficiency in reasoning datasets at the multilingual level.
Multi-Language Code Benchmark. Multilingual evaluation datasets are important for assessing the comprehensive coding capabilities of code LLMs. In the early stages, multilingual code datasets were mainly used for code translation tasks (Elnaggar et al., 2021; Ahmad et al., 2021; Roziere et al., 2020, 2021; Zhu et al., 2022b; Yan et al., 2023; Zhu et al., 2022a). These datasets often consist of problem solutions in different languages extracted from algorithm competition-related websites, thus suffering from data contamination issue. Bechmark like McEval (Chai et al., 2024), which relies on human annotation, requires a high cost. In this work, we provide a process using LLMs for multilingual code translation, which can achieve a high accuracy and low cost in creating a multilingual dataset.
6 Conclusion
@@
7 Limitations
There are Three limitations in this work. First, our benchmark relies on model-generated data. While this approach effectively reduces costs and avoids issues of data leakage, it may introduce biases in the dataset towards the LLMs that generated the data. However, we select different LLMs to construct evaluation data and conduct experiments on this data. The experimental results indicate that our dataset construction method does not affect the fairness of the dataset. Additionally, the model-based translation method cannot guarantee that all data will be perfectly translated into other PLs. However, through our method, based on 800 pieces of data, we generated 500 aligned pieces of data for 19 PLs. From our evaluation results, it can be seen that this amount of data is sufficient to effectively distinguish the reasoning capabilities of each LLM. Finally, translated codes from Python cannot reflect several specific language features of other languages. Our intention is to achieve highly aligned multilingual data across 19 languages. However, preserving language-specific features make it challenging to create such aligned data, as the distinctive features of one language are inherently difficult to replicate in others. This represents a trade-off. In our future work, we plan to explore how to construct multilingual data while retaining more language-specific features.
Appendix A Dataset Bias for Data-Generating LLM
The experimental results of existing papers have already demonstrated that the LLMs used for generate CRUXEVAL benchmark do not introduce unfairness Gu et al. (2024). Therefore, our goal is to verify whether our pipeline can lead to unfairness on datasets of other programming languages (PLs). To achieve this, we reduce the number of iterations and use a single model for data generation, aiming to determine if different models can gain an advantage on the data they generated themselves.
We select Qwen2.5-Coder-32B-Instruct (Qwen 32B), deepseek-coder-33b-instruct (Dpk 33B), and deepseek-coder-6.7b-instruct (Dpk 7B) to generate data respectively, and reduce the maximum number of iterations to 5. We generated data for both commonly used PLs C++ and less commonly used PLs Racket. After taking the intersection, we obtain 566 aligned C++ data points and 206 aligned Racket data points. We then calculate the pass@1 for these three models, with the specific results shown in Table 3.
Input Reasoning (C++)
Data Evaluate Qwen 32B Dpk 33B Dpk 7B
Qwen 32B 76.3 49.6 38.3
Dpk 33B 75.6 47.5 39.2
Dpk 7B 76.3 48 37.6
Output Reasoning (C++)
Data Evaluate Qwen 32B Dpk 33B Dpk 7B
Qwen 32B 65.5 53.0 42.6
Dpk 33B 67.7 51.2 44.3
Dpk 7B 67.1 52.6 43.2
Input Reasoning (Racket)
Data Evaluate Qwen 32B Dpk 33B Dpk 7B
Qwen 32B 82.0 51.9 42.7
Dpk 33B 77.2 49.0 37.9
Dpk 7B 78.6 42.2 41.3
Output Reasoning (Racket)
Data Evaluate Qwen 32B Dpk 33B Dpk 7B
Qwen 32B 78.6 58.7 47.6
Dpk 33B 76.2 57.8 46.1
Dpk 7B 79.1 56.8 44.7
Table 3: Impact of data generating LLMs
It can be observed that the relative ordering of the three models does not change. Moreover, the model used for data generation does not significantly impact the final evaluation results. For instance, in the Input Reasoning task for C++, Qwen 32B’s pass@1 consistently remain around 76.
We analyze the possible reason: During both the data generation phase and the evaluation phase, the tasks performed by the LLMs are different. We use the model to generate code in different languages (i.e., code translation), while we evaluate the model’s ability to infer outputs/inputs based on given inputs/outputs in code (i.e., code reasoning). It is important to note that the test cases used during the code reasoning phase are not visible to the model during the code translation phase, ensuring a fair evaluation process. Therefore, the model is performing two distinct tasks (i.e., code translation and code reasoning), and there is no unfairness involved.
Appendix B The Improvements of Pipline
B.1 enhance of the pipline
@@ Appendix B The Improvements of Pipline
B.2 transform complex types
@@
Listing 4 An example with complex type
@@
Appendix C The Difficulties for Translating
(1) For indexing functions, the starting position is not 0 but 1, with Julia being a typical language that exhibits this behavior. An example is shown in listing 5. (2) For the conversion between Python’s str type and its own char and string types, D language is a typical case where this issue arises. An example is shown in listing 6. (3) For the transformation and comparison of dictionary types. For example, C# need to first sort two Dictionary before compare them.
Listing 5 Julia Error Case
1function f(text::String, s::Int64, e::Int64)::Int64
2 sublist = text[s+1:e] # Julia uses 1-based indexing, so adjust the start index
3 if isempty(sublist)
4 return -1
5 end
6 min_char = minimum(sublist)
7 return findfirst(==(min_char), sublist) - 1 # Adjust for 0-based index in the result
8end
Appendix D Details of Benchmark Construction
D.1 experimental setup of each step
In Generation step. Given the higher usage cost of closed-source LLMs, we set $N$ to $5$, $k$ to $5$, $\delta A$ to 00, and the temperature to 0.2 for GPT3.5-Turbo. For DeepseekCoder-33B-Instruct We set $N$ to $50$, $k$ to $5$, $\delta A$ to 00, and the temperature to 0.8. In Repair step. We set temperature to 0. For GPT-4o, we generate once, and correct errors three times.
D.2 prompt of each step
In Figures 7 and 6, we include the prompts we use for our benchmark construction. We use a few-shot prompt for all models. For Generation step of each model, the prompt is show in Figure 7. The example in this Figure is used for GPT3.5-turbo. After the Generation and Repair of GPT3.5-turbo, we choose three examples from the correct generated problems which include str, List, Dict type respectively. We use these examples for Deepseekcoder-instruct-33b Generation. For GPT-4o, based on the summarized difficulties in translation, we provide three examples, as shown in Listing 6, 7, and 8. we present these examples in three distinct languages. For each specific language translation, we employ the corresponding language version of the three examples.